The Curly (Cy) mutant in fruit flies causes curled up wings. This mutant is due to an autosomal dominant mutation that exhibits 80% penetrance. The miniature body (m) phenotype is due to an X-linked recessive mutation and is completely penetrant. A male with normal body size and heterozygous for the Cy mutation is crossed to a female with normal wings and miniature body. What is the probability of obtaining progeny flies from this cross that are males with miniature bodies and curly wings?

A. 1/4

B. 1/5

C. 1/10

D. 1/6

E. 0

Please show work
Answer #1

The answer will be 1/5 (Option B).

Explanation: Let, Cy+ = normal wing allele, Cy = Mutant curly wing allele, Xm = X chromosome with miniature body mutation, X = normal X chromosome

We will differently find probability of males with miniature bodies & curly wings from a cross between male with normal body size & heterozygous for the Cy mutation & a female with normal wings & miniature body. At last we will combined this probability values to determine our result.

Female with miniature body will have genotype XmXm (As miniature body is an X-linked recessive mutation) & male with normal body size will have genotype XY. The cross between them is shown in below Punnett square.

Gametes Xm Xm
X XXm (Normal female) XXm (Normal female)
Y XmY (Male with miniature body) XmY (Male with miniature body)

So, probability of male with miniature body from this cross is 2/4 or 1/2.

Male with heterozygous for Cy mutation will have genotype Cy+Cy & female with normal wings has genotype Cy+Cy+. The cross between them is shown in below Punnett square.

Gametes Cy+ Cy+
Cy+ Cy+Cy+ (Normal wing) Cy+Cy+ (Normal wing)
Cy Cy+Cy (Curled wing) Cy+Cy (Curled wing)

So, probability of curled wing = 2/4 x 0.8 (As mutation penetrance is 80%) = 0.4

So, probability of obtaining males with miniature bodies & curly wings = Probability of male with miniature body x Probability of curled wing = 1/2 x 0.4 = 1/5